if
w/x =6/11
y/z=16/23 and
w/y=9/16
then what is the valve of x/y?

Answer by richwmiller(17219)

(w/x) =6/11, (y/z)=16/23,(w/y)=9/16
(w/y)/(w/x)=x/y
(9/16)/(6/11)=33/32=x/y